OpenStudy (anonymous):

Ok so Can someone walk me through this? System of Equations 1. Y - 3x -11 Y = -4x +7

OpenStudy (anonymous):

Where is the equal sign in the first equation?

OpenStudy (danjs):

if y=something and y=something else then something = something else 3x-11=-4x+7 solve for x, use that x value to solve either equation for y

OpenStudy (anonymous):

Ok let me solve it. :) thank you

OpenStudy (anonymous):

hows the 3x positive?

OpenStudy (jackthegreatest):

u plug it in

OpenStudy (jackthegreatest):

since y= 3x-11, then u can just plug in that y value into the equation @Keridavis

OpenStudy (anonymous):

my teacher taught it different.. so so different... lordd!

OpenStudy (jackthegreatest):

well there is another way

OpenStudy (jackthegreatest):

u can do this

OpenStudy (jackthegreatest):

u can do the subtraction method Y = 3x -11 y = -4x +7 ----------- 0 = 7x -18 -7x = -18 7x=18 @Keridavis

OpenStudy (jackthegreatest):

can u solve the rest?

OpenStudy (jackthegreatest):

after u solve for x, then plug it into the equation and u can get y

OpenStudy (jackthegreatest):

can i have a medal @Keridavis

OpenStudy (anonymous):

eh...